# Graph Language Models

### v1

@@ Abstract
While Language Models have become workhorses for NLP, their interplay with textual knowledge graphs (KGs) – structured memories of general or domain knowledge – is actively researched. Current embedding methodologies for such graphs typically either (i) linearize graphs for embedding them using sequential Language Models (LMs), which underutilize structural information, or (ii) use Graph Neural Networks (GNNs) to preserve graph structure, while GNNs cannot represent textual features as well as a pre-trained LM could. In this work we introduce a novel language model, the Graph Language Model (GLM), that integrates the strengths of both approaches, while mitigating their weaknesses. The GLM parameters are initialized from a pretrained LM, to facilitate nuanced understanding of individual concepts and triplets. Simultaneously, its architectural design incorporates graph biases, thereby promoting effective knowledge distribution within the graph. Empirical evaluations on relation classification tasks on ConceptNet subgraphs reveal that GLM embeddings surpass both LM- and GNN-based baselines in supervised and zero-shot settings. [^1] https://github.com/Heidelberg-NLP/GraphLanguageModels
@@ 1 Introduction
Knowledge Graphs (KGs) play a crucial role in information representation and retrieval, offering a structured framework to organize and connect vast data. KGs excel in making manifold relationships explicit, empowering organizations and researchers to unlock hidden insights and enhanced information retrieval for improved decision-making. As the data landscape expands, KGs become essential for navigating and harnessing the wealth of information in the digital age, enabling the discovery of unseen connections within and across complex data sources, ultimately driving innovation.
Language models (LMs) have demonstrated remarkable proficiency in various NLP tasks, showcasing their capabilities in analyzing and generating text. However, they are predominantly confined to processing sequential data, such as sentences or paragraphs, which limits their ability to effectively handle graph-based data. I.e., graphs – consisting of nodes and edges that represent relationships and interactions in a non-linear way – present a challenge that goes beyond the sequential structure of natural language.
@@
Many KGs consist of a collection of knowledge triplets, where nodes are entities and edges represent relations holding between these entities. Each triplet represents a fact, e.g., (dog; IsA; animal) in ConceptNet (Speer et al., 2017) or (Thailand; Capital; Bankok) in DBpedia (Auer et al., 2007). Despite the (usual) simplicity of each triplet individually, complex structures emerge when triplets are combined in a KG. Such Graphs of Triplets (GoT) are common representations for KGs.
@@
For effectively using GoTs, we need meaningful encodings of their components. A natural choice is leveraging LMs, as they can capture the semantic content of textually encoded entities, relations or entire triplets. However, LMs are not inherently prepared to capture graph-structured information, which is crucial to capture interactions in a GoT. To alleviate this problem, we can leverage graph NNs (GNNs). GNNs, on the other hand, are not well prepared to capture meanings associated with text, and hence often LMs are used to convert nodes (and potentially edges) to language-based semantic embeddings. But in such settings, semantic encoding leveraged from LMs and structural reasoning performed by GNNs are separated and are driven by distinct underlying principles. We expect this to limit model performance if both, semantic and structural information, are important for a task.
@@
To enable early fusion of semantic and structural information, we propose a new Graph Language Model (GLM) formulation. The most powerful LMs today are transformer-based. Since transformers operate on sets, so-called Positional Encoding (PE) is used to make them sensitive to the inherent sequential ordering of linguistic inputs. In our GLM formulation, we modify PE and self-attention to convert LMs (i.e., sequence transformers) to graph transformers that natively operate on graphs, while preserving their LM capabilities. Usually, a new model architecture requires pretraining from scratch, which is extremely costly. By creating a new architecture that unifies relevant features of the LM and Graph Transformer (GT) architectures, we can continue using the pretrained parameters of the LM. This is possible since we adopt some non-invasive changes in the self-attention module: We define LM-like attention patters for the processing of linearly organized textual information, and GT-like attention patterns for information aggregation in the graph. Hence, our proposed GLM inherits semantic understanding of entities and triplets from the LM, while architectural features assimilated from the GT allow it to directly perform structural reasoning, without relying on external GNN layers.
@@
Our contributions are: (i) We propose Graph Language Models (GLMs) and offer a theoretical framework to construct them. GLMs are graph transformers, which enables graph reasoning. Simultaneously, they integrate elements of the transformer LM architecture, which allows them to effectively represent and contextualize triplets inside a GoT. (ii) Experiments on relation classification in ConceptNet subgraphs show that GLMs outperform LM- and GNN-based methods for encoding GoTs – even when LM parameters are not updated to adapt to the new architecture in the GLM. (iii) Ablations confirm that weight initialization from a LM is crucial for the GLM’s high performance.
2 Preliminary: Graph Transformers
@@
This section gives a brief overview of graph transformers (GTs), focusing on architectural choices relevant for our work. We refer to Min et al. (2022) for a detailed overview of GTs. As GTs are a type of GNN, we also discuss some general properties of GNNs that motivate our design choices in §3.
The attention mechanism in self-attention can be written as
@@
where $Q$, $K$ and $V$ are the query, key and value matrices, and $d$ is the query and key dimension (Vaswani et al., 2017). $B_{P}$ and $M$ are matrices which can be used for positional encoding and masking, respectively. Setting $B_{P}=M=0$ yields the formulation in Vaswani et al. (2017). We will discuss the roles of $B_{P}$ and $M$ for GTs below.
@@
The Self-Attention mechanism of transformer models is permutation invariant, i.e., it doesn’t have any notion of the order of elements in its input. Positional Encoding (PE) is commonly used to inform LMs on the ordering of tokens in a sentence (Dufter et al., 2022). Most approaches employ either absolute PE, where absolute token positions are encoded (Vaswani et al., 2017; Gehring et al., 2017) or relative PE, which encodes the relative position between pairs of tokens (Shaw et al., 2018; Raffel et al., 2020; Su et al., 2021; Press et al., 2022). Absolute PE are typically combined with the input sequence and hence, the PE does not need to be encoded in self-attention ($B_{P}=0$). In relative PE $B_{P}$ is used to encode a bias depending on the relative distances between pairs of tokens – for example, by learning one scalar for each possible distance:
@@
where $P$ is a matrix of relative distances and $f(\cdot)$ an elementwise function. Some approaches are orthogonal to this classification and can be combined with either one (Ke et al., 2021; Ruoss et al., 2023).
Similarly, GTs use PEs to encode the structure of the input, and hence, their PE has to encode a graph structure, as opposed to a sequence. This can again be done with absolute or relative PEs. However, defining an “absolute position” of a node or edge in a graph is not straightforward. While many methods exist, they are not directly compatible with the usual (absolute) “counting position” known from sequence encoding in LMs. We therefore focus on relative PE in this work. Given a directed acyclic path in a graph, we can define the (signed) distance between any pair of nodes along a path simply as the number of hops between the nodes. The sign can be determined by the direction of the path. Thus, when we find a consistent set of such paths, we have PE for our graph. Our choice of these paths will be described in §3.
@@
On the other hand, it has been shown that a global view of the graph can enable efficient, long-ranged information flow (Alon and Yahav, 2021; Ribeiro et al., 2020). We will therefore present two model variants – a local and a global GLM.
(a) Original GoT
(b) Extended Levi graph of GoT (with relative distances $P$ for dog)
@@ 3 Graph Language Model
We aim to design an architecture that can efficiently and jointly reason over textual and graph-structured data. GTs are a special type of GNN (cf. §2 and Bronstein et al. (2021)), and hence can offer the desired graph priors. Yet, they lack language understanding, as they are not equipped with knowledge that LMs acquire during pretraining on textual data.
One intuitive approach allowing to bridge this gap is to pretrain a GT from scratch (Schmitt et al., 2021). But large-scale pretraining is costly and the necessary data is bound to be scarce. We therefore take a different avenue. We hypothesize that the language understanding capabilities required for reasoning over GoTs is similar to the reasoning performed when processing sequential texts. Intuitively this should be the case, since (i) GoTs are designed to be understandable by humans and (ii) literate people can ‘read’ and understand GoTs. Existing LMs have such language understanding capability stored in their parameters.
By initializing a GT with parameters from a compatible LM, we obtain our Graph Language Model (GLM). The GT architecture introduces graph priors, while the parameter initialization from the LM introduces textual understanding. In the following we explain the necessary modifications to the input graph and to the model, to make this work. The general idea is that (verbalized) triplets should resemble natural language as much as possible to enable LM weights to capture them, while graph reasoning should be done via message passing.
@@
A LM tokenizer converts text into a sequence of tokens, which the LM can then process. In a similar way, we need to process GoTs, such that the GLM can process the graphs “as a LM would do” (cf. Fig. 2). To achieve this, we first convert the GoT to its so-called Levi graph (Schmitt et al., 2021), i.e., we replace each edge with a node that contains the relation name as textual feature, and connect the new node to the head and tail of the original edge via unlabeled edges, preserving the directionality of the original edge. If necessary, we verbalize the relation: for example, for ConceptNet (CN) we convert the relation label IsA to the natural language expression is a. Next, we tokenize each node using the tokenizer of the underlying LM. Finally, we split each node into multiple nodes, such that every new node contains a single token, and connect adjacent nodes via edges, again preserving the directionality of the original relation. This yields the final extended Levi graph (see Fig. 1(b)).
In this representation, each triplet is formalized as a sequence of tokens – just as it would be for a standard LM. The representation is thus well-suited for the GLM to seamlessly process of the original graph nodes and labels. [^2] Note, however, that the token sequence of the converted GoT is not necessarily perfectly identical to the sequence of tokens that corresponds to the input triplets. We tokenize each node in the Levi graph individually, to ensure consistent tokenization of concepts shared by multiple triplets. This removes whitespace between concepts and relations, which impacts tokenization. We leave it for future work to investigate the impact of this effect.
Figure 3: Relative position matrix $P$ for tokens in Fig. 1(b). Entries with G2G either have no relative position ($\ell$GLM) or a relative position of $\infty$ ($g$GLM) (cf. Appendix, Fig. 4).
@@
As discussed in §2, we prefer PEs that encode the relative position between pairs of tokens, determined by their signed distance. We can directly adopt this method to encode the relative position of pairs of tokens occurring within the same triplet – by simply considering the triplet as a piece of text, and counting the token distance in this text. Note that a single token can occur in multiple triplets, having, e.g., multiple “lefthand side neighbors” (cf. animal in Fig. 1(b) and 3). While this does not occur in ordinary sequential text, it does not impose a problem for relative PE.
Yet, for pairs of tokens that do not belong to the same triplet, we cannot follow the usual approach, since we cannot guarantee that there is a canonical text that connects the tokens. To determine the distance between such pairs of tokens belonging to different triplets, previous work considered, e.g., the length of the shortest path between them (Schmitt et al., 2021). However, this results in PEs that do not come natural to a LM, since a triplet would appear in reversed order, if a triplet is traversed in “the wrong direction” in the shortest path. [^3] For example, cat would see the graph as the following sequence: cat is a animal a is dog a is poodle black. We therefore omit structure-informed PE between tokens that do not belong to the same triplet and instead propose two GLM variants: a local ($\ell$GLM) and a global ($g$GLM) Graph Language Model.
Metric $r=1$ $r=2$ $r=3$ $r=4$ $r=5$
Table 1: Structural statistics of train graphs.
@@
However, it has been shown that a global view can sometimes be beneficial (Ribeiro et al., 2020). Hence, we also formalize the $g$GLM, as an alternative, where self-attention can connect any node to every other node. For this setting we need to assign a PE to any pair of tokens, including those that do not occur within the same triplet. For these pairs we set the relative position to +$\infty$+ ∞ in $P$. In a LM a relative PE of +$\infty$+ ∞ means that the respective tokens occur somewhere “far” away in a remote text passage. LMs learn a proximity bias during pretraining, i.e., they tend to have higher attention scores between tokens that are close to each other in the text. This means that tokens with a high relative distance tend to have low attention scores. For our $g$GLM this corresponds to a graph bias where distant nodes are less important, but are still accessible. Note that unlike in the $\ell$GLM, this bias is not part of the architecture. Instead, it originates from the pretrained parameters, which means that the $g$GLM can learn to attend more to distant tokens.
Together with $P$ and $M$, the GLM takes a sequence consisting of all tokens in the extended Levi graph as input. For this, we technically need to “linearize” the graph. However, the order of tokens in the resulting sequence does not matter: relative positions in $P$ are determined by distances in the the graph, not in the sequence. Permuting the input sequence simply means that rows and columns of $P$ and $M$ need to be permuted accordingly, but the resulting token embeddings remain unchanged. Example matrices for $P$ and $M$ for the $\ell$GLM and $g$GLM are shown in Appendix, Fig. 4.
@@
If the LM’s self-attention is uni-directional, information can only be propagated in the direction of the arrows in Fig. 1(b) for the $\ell$GLM. This means that, e.g., the representation of the node black poodle is independent of the rest of the graph. In principle, we can augment the graph with inverse relations, to enable bidirectional information flow even with a unidirectional LM. In this work, however, we restrict our analysis to bidirectional models.
@@
In this work we chose T5 (Raffel et al., 2020) as base LM to instantiate the local and global GLM variants. T5 comes in various sizes and consists of a bidirectional encoder, together with a unidirectional decoder. We convert the encoder to a GLM using the steps above. In T5, the relative distances in $P$ are grouped into so-called buckets, and each bucket gets mapped to one learned positional bias in $B_{P}$ for each head. The positional biases are shared across layers. The T5 decoder is not needed to encode graphs, but could be used to generate a sequence, e.g., text or a linearized graph.

### v3

@@ Abstract
While Language Models (LMs) are the workhorses of NLP, their interplay with structured knowledge graphs (KGs) is still actively researched. Current methods for encoding such graphs typically either (i) linearize them for embedding with LMs – which underutilize structural information, or (ii) use Graph Neural Networks (GNNs) to preserve the graph structure – but GNNs cannot represent text features as well as pretrained LMs. In our work we introduce a novel LM type, the Graph Language Model (GLM), that integrates the strengths of both approaches and mitigates their weaknesses. The GLM parameters are initialized from a pretrained LM to enhance understanding of individual graph concepts and triplets. Simultaneously, we design the GLM’s architecture to incorporate graph biases, thereby promoting effective knowledge distribution within the graph. This enables GLMs to process graphs, texts, and interleaved inputs of both. Empirical evaluations on relation classification tasks show that GLM embeddings surpass both LM- and GNN-based baselines in supervised and zero-shot setting, demonstrating their versatility. [^1] https://github.com/Heidelberg-NLP/GraphLanguageModels
@@ 1 Introduction
Knowledge Graphs (KGs) are essential for organizing vast data, to facilitate information retrieval, or revealing hidden insights for decision-making (Plenz et al., 2024). KGs excel in explicitly representing manifold relationships, so with an expanding wealth of information they become crucial tools in the digital age.
@@
Many KGs consist of knowledge triplets, where nodes are entities and edges represent relationships holding between them. Each triplet represents a fact in pseudo-natural language, e.g., (Thailand; Capital; Bangkok) in DBpedia (Auer et al., 2007). Despite the (usual) simplicity of each individual triplet, complex structures emerge in KGs. We refer to such KGs as Graphs of Triplets (GoTs).
@@
To use GoTs effectively, we need meaningful encodings of their components. A natural choice is leveraging LMs, as they can capture the semantics of textually encoded entities, relations or entire triplets. But LMs are not prepared to capture graph-structured information and cannot model complex interactions in a GoT. To alleviate this problem, one can leverage graph NNs (GNNs). But GNNs are not well suited to capture meanings associated with text, and hence often LMs are used to convert nodes (and possibly edges) to language-based semantic embeddings. But in such settings, semantic encoding leveraged from LMs and structural reasoning performed by GNNs are separated and are driven by distinct underlying principles. We expect this to limit model performance if both textual and structural information are important for a task.
@@
In this work we introduce a Graph Language Model (GLM) that resolves this tension through early fusion of textual and structural information. Most LMs today are transformers. Since transformers operate on sets, Positional Encoding (PE) is used to inform them about the inherent sequential ordering of linguistic inputs. In our GLM formulation, we modify PE and self-attention to convert LMs (i.e., sequence transformers) to graph transformers that natively operate on graphs, while preserving their LM capabilities. Usually, a new architecture requires pretraining from scratch, which is extremely costly. By adopting some non-invasive changes in the LM’s self-attention module, we transform the LM to a Graph Transformer (GT) – while maintaining compatibility with its pretrained LM parameters. When encoding a graph, LM-like attention patterns process linearly organized textual information of individual triplets, while GT-like attention patterns aggregate information along the graph structure. Hence, the GLM inherits text understanding of triplets from the LM, while its GT architecture allows it to directly perform structural reasoning, without additional GNN layers.
Importantly, for text sequences – which can be seen as a special type of graph – the GLM is identical to the original LM. This allows the GLM to process interleaved inputs of text and GoT jointly, handling both modalities in a single framework.
@@
Our main contributions are: (i) We propose Graph Language Models (GLMs) and a theoretical framework to construct them. GLMs are graph transformers, which enables graph reasoning. Simultaneously, they inherit and exploit LM weights, enabling them to represent and contextualize triplets in a GoT. Further, by encoding texts and graph components alike, it can naturally take graph and text data as interleaved inputs. (ii) Experiments on relation classification in ConceptNet subgraphs show that GLMs outperform LM- and GNN-based methods for encoding GoTs – even when the inherited LM parameters are not updated during GLM training. (iii) KG population experiments on Wikidata subgraphs and corresponding Wikipedia abstracts show that GLMs can reason over interleaved inputs of GoTs and text – again, outperforming strong LM-based methods.
2 Related Work
LMs
One way to augment LMs with knowledge from KGs (Pan et al., 2024) is to formulate pretraining objectives that operate on a KG. E.g., LMs can be trained to generate parts of a KG, encouraging the LM to store KG content in its parameters. Typically, single triplets are used for pretraining (Bosselut et al., 2019; Wang et al., 2021; Hwang et al., 2021; West et al., 2023). In such cases, the graph structure is not a target of pretraining. Some works generate larger substructures, such as paths or linearized subgraphs (Wang et al., 2020a; Schmitt et al., 2020; Huguet Cabot and Navigli, 2021). In either case, the LM needs to memorize the KG, as it will not be part of the input during inference.
Another approach is to provide the linearized KG as part of the input during inference. This is common for KG-to-text generation (Schmitt et al., 2020; Ribeiro et al., 2021; Li et al., 2021), where models learn to take the linearized (typically small-sized) graphs as input. A recent trend is retrieval augmented generation, where relevant parts of a knowledge base, or KG, are retrieved, linearized and provided as part of a prompt (Gao et al., 2024). [^2] Cf. www.llamaindex.ai and www.langchain.com
In both options the graph must be linearized to fit the input or output of a sequence-to-sequence LM. Hence, no graph priors can be enforced – instead, the LM has to learn the graph structure implicitly. By contrast, GLMs model a graph as a true graph and have inductive graph priors instilled in their architecture. This prepares a GLM for more proficient graph reasoning, compared to a LM approach.
GNNs
LMs excel at representing single triplets, but struggle with structural reasoning. To alleviate this problem, LMs can be combined with GNNs. Many approaches get node and edge features from LMs and aggregate this information in the graph with GNNs (Lin et al., 2019; Malaviya et al., 2020; Zhao et al., 2023). Zhang et al. (2022); Yasunaga et al. (2022) train models consisting of a LM and a GNN that encode interleaved text and graph inputs jointly. They also use a LM to obtain node features.
While some approaches jointly train for textual understanding and graph reasoning, none offer a unified method. By contrast, our GLM formulation seamlessly integrates both in a holistic framework for embedding language and KGs.
Graph Transformers
GTs, a special type of GNN (Bronstein et al., 2021), gain popularity in NLP and beyond (Min et al., 2022; Müller et al., 2023). E.g., Koncel-Kedziorski et al. (2019) and Wang et al. (2020b) train GTs to generate text from KGs and AMRs, respectively. Most relevant to our work is Schmitt et al. (2021), who use GTs for KG-to-text generation. Similar to us, they employ PE matrices, but train their model from scratch, which limits its applicability: while their model trained on WebNLG (Gardent et al., 2017) has a vocabulary size of 2,100, initializing a GLM from T5, equips it with T5’s full vocabulary of 32,128 tokens.
Concurrently, and independently from our work, Li et al. (2024) also convert a LM to a graph transformer. They focus on data-to-text generation, where they unify table, key-value and KG structures in a unified graph format, and apply structure-enhanced pre-training to support data-to-text generation with their structure-enhanced transformer model. They apply attention maps similar to ours to better capture the graph-structured input, which the pre-trained model rewrites into natural language. Contrary to their work, we do not resort to structure-enhanced pre-training – which is restricted in resources – but instead assess the GLMs’ innate capabilities. We showcase the versatility of the inherited LM parameters in conjunction with our graph transformer architecture, by applying them to challenging reasoning tasks, where the model needs to reason over complementary inputs from text and graphs, and where it needs to infer information not present in the input, unlike data-to-text generation. Moreover, we demonstrate that our architectural changes are highly compatible with the original LM weights, via linear probing experiments, where the GLM outperforms conventional LM and Graph Transformer models.
3 Preliminary: Graph Transformers (GT)
@@
This section briefly introduces graph transformers, focusing on architectural choices relevant for our work. We also discuss some general properties of GNNs that motivate our design choices in §4.
The attention in self-attention can be written as
@@
where $Q$, $K$ and $V$ are the query, key and value matrices, and $d$ is the query and key dimension. The $B_{P}$ and $M$ matrices can be used for positional encoding and masking. Setting $B_{P}=M=0$ yields the standard formulation (Vaswani et al., 2017).
@@
The self-attention mechanism of transformer models is permutation invariant, i.e., it doesn’t have any notion of the order of its input elements. Thus, positional Encoding (PE) is used to inform LMs of the ordering of tokens in a text (Dufter et al., 2022). Most approaches employ either absolute PE, where absolute token positions are encoded (Vaswani et al., 2017; Gehring et al., 2017) or relative PE, which encodes the relative position between pairs of tokens (Shaw et al., 2018; Raffel et al., 2020; Su et al., 2021; Press et al., 2022). Absolute PE is typically combined with the input sequence and hence, the PE does not need to be encoded in self-attention ($B_{P}=0$). For relative PE, $B_{P}$ encodes a bias depending on the relative distances between pairs of tokens – for example, by learning one scalar for each possible distance:
@@
where $P$ is a matrix of relative distances and $f(\cdot)$ an elementwise function.
Similarly, GTs use PEs to encode the structure of the input, and hence, their PE has to encode a graph structure, as opposed to a sequence. This can again be done with absolute or relative PEs. However, defining an “absolute position” of a node or edge in a graph is not straightforward. While many methods exist, they are not directly compatible with the usual (absolute) “counting position” known from sequence encoding in LMs. In this work we thus focus on relative PE. Given a directed acyclic path in a graph, we can define the (signed) distance between any pair of nodes along a path simply as the number of hops between the nodes. The sign can be set by the direction of the path. Thus, by finding a consistent set of such paths in §4, we obtain relative distances and hence the graph’s PE.
@@
On the other hand, it has been shown that a global view of the graph can enable efficient, long-ranged information flow (Alon and Yahav, 2021; Ribeiro et al., 2020). We will therefore present two model variants in §4 – a local and a global GLM.
(a) Original GoT.
(b) Extended Levi graph of GoT (with relative distances $P$ for dog).
@@ 4 Graph Language Model
We aim to design an architecture that can efficiently and jointly reason over text and graph-structured data. GTs can offer desired graph priors, but they lack language understanding.
One intuitive approach to bridge this gap is to pretrain a GT from scratch (Schmitt et al., 2021). But pretraining is costly and the necessary data bound to be scarce. We thus take a different avenue. We hypothesize that for reasoning over GoTs a model needs language understanding capabilities similar to those used for reasoning over text. Intuitively this should be the case, since (i) GoTs are designed to be understandable by humans and (ii) literate people can “read” and understand GoTs.
By initializing a GT with parameters from a compatible LM, we obtain our Graph Language Model (GLM). The GT architecture introduces graph priors, while parameter initialization from the LM gives it language understanding capabilities. In the following we explain the necessary modifications to the input graph and to the model, to make this work. The general idea is that (verbalized) triplets should resemble natural language as much as possible to enable LM weights to capture them, while graph reasoning should work via message passing.
@@
A LM tokenizer converts text into a sequence of tokens from the LM vocabulary. Similarly, we process GoTs, such that the GLM can process the graphs “as a LM would do” (cf. Fig. 2). To achieve this, we first convert the GoT to its so-called Levi graph (Schmitt et al., 2021), i.e., we replace each edge with a node that contains the relation name as text feature, and connect the new node to the head and tail of the original edge via unlabeled edges, preserving the direction of the original edge. Next, we tokenize each node and split each node into multiple nodes, such that every new node corresponds to a single token. New edges connect adjacent nodes, again preserving the original direction. This yields the extended Levi graph (see Fig. 2(b)). In this representation, each triplet is represented as a sequence of tokens – just as it would be for a standard LM. [^3] Note that the token sequence of the converted GoT is not necessarily perfectly identical to the token sequence that corresponds to the input triplets. We tokenize each node in the Levi graph individually, to ensure consistent tokenization of concepts shared by multiple triplets. This removes whitespace between concepts and edges, which impacts tokenization. We leave investigation of the impact of this effect to future work.
Figure 3: Relative position matrix $P$ for tokens in Fig. 2(b). Entries with G2G have no relative position ($\ell$GLM) or are initialized from +$\infty$+ ∞ ($g$GLM). Cf. §A.
@@
As discussed in §3, we prefer PEs that encode the relative position between pairs of tokens, determined by their signed distance. We can directly adopt this method to encode the relative position between pairs of tokens occurring within the same triplet – by simply considering the triplet as a piece of text, and counting the token distance in this text. Note that a single token can occur in multiple triplets, leading to, e.g., multiple “left-hand side neighbors” (cf. animal in Fig. 2(b) and 3). While this does not occur in ordinary sequential text, it does not impose a problem for relative PE.
Yet, the approach above breaks when tokens do not belong to the same triplet. To determine the distance between such pairs of tokens, previous work considered, e.g., the length of the shortest path between them (Schmitt et al., 2021). However, this results in PEs that do not come natural to a LM, since a triplet would appear in reversed order, if it is traversed in “the wrong direction” in the shortest path. [^4] For example, cat would see the graph as the following sequence: cat is a animal a is dog a is poodle black. We therefore omit structure-informed PE between tokens that do not belong to the same triplet and instead propose two GLM variants: a local ($\ell$GLM) and a global ($g$GLM) one.
@@
However, it has been shown that a global view can have benefits (Ribeiro et al., 2020). Hence, we also formalize the $g$GLM, as an alternative where self-attention can connect any node to every other node. For this setting we need to assign a PE to any pair of tokens, including those that do not occur within the same triplet. For these pairs we introduce a new graph-to-graph (G2G) relative position. LMs don’t have learned parameters for G2G connections, so we initialize the parameters with the corresponding parameters of a relative position of +$\infty$+ ∞. In a LM a relative position of +$\infty$+ ∞ means that the respective tokens occur somewhere “far” away in a remote text passage. LMs learn a proximity bias during pretraining, i.e., they tend to have higher attention scores between tokens that are close to each other in the text. This means that tokens with a high relative distance tend to have low attention scores. For our $g$GLM this corresponds to a graph bias where distant nodes are less important, but are still accessible. [^5] Preliminary experiments showed that initializing G2G parameters from +$\infty$+ ∞ outperforms random initialization, which outperforms initialization from 00. Note that unlike in the $\ell$GLM, this bias is not part of the architecture. It originates from the pretrained parameters, meaning that the $g$GLM can learn to attend to distant tokens.
Along with $P$ and $M$, the GLM takes a sequence of all tokens in the extended Levi graph as input. For this, we technically need to “linearize” the graph. However, the order of tokens in the resulting sequence does not matter: relative positions in $P$ are determined by distances in the graph, not in the sequence. Permuting the input sequence simply means that rows and columns of $P$ and $M$ need to be permuted accordingly, but the resulting token embeddings remain unchanged. See example matrices for $P$ and $M$ for $\ell$GLM and $g$GLM in §A.
Being transformers, GLMs have the same computational complexity as their respective LM. For sparse graphs the $\ell$GLM could make use of sparse matrix multiplication, making it more efficient than a corresponding LM or $g$GLM. However, for our experiments this was not necessary.
Joint graph and text encoding
If we use normal matrices for $P$ and $M$, the GLM is identical to its underlying LM. Hence, GLMs can be applied to texts and – more interestingly – interleaved inputs of text and graph. In this joint setting, $P$ and $M$ each consists of four sub-matrices that correspond to self-attention between tokens from (i) graph-to-graph, (ii) text-to-text, (iii) text-to-graph and (iv) graph-to-text. Graph-to-graph sub-matrices are formatted as described above for $\ell$GLM and $g$GLM, respectively. Text-to-text sub-matrices are standard matrices from conventional sequence transformers. We introduce new T2G and G2T relative positions for text-to-graph, and graph-to-text connections, respectively. With this, the model can learn interaction strength between the two modalities. Similar to G2G in $g$GLM, we initialize T2G and G2T parameters from +$\infty$+ ∞. See example matrices in §A.
@@
If a LM’s self-attention is unidirectional, information can only propagate along the direction of arrows in Fig. 2(b) for the $\ell$GLM. This means that, e.g., the representation of the node black poodle is independent of the rest of the graph. We could augment the graph with inverse relations to enable bidirectional information flow with unidirectional LMs, but in this work, we restrict our analysis to bidirectional models.
@@
We use T5 (Raffel et al., 2020) – a bidirectional encoder with unidirectional decoder – as base LM to instantiate GLMs. In T5, relative distances in $P$ group into so-called buckets, and each bucket maps to one learned positional bias in $B_{P}$ for each head. Positional biases are shared across layers. The decoder is not needed to encode graphs, but can be used to generate sequences, such as text or linearized graphs in future work.